λ for one line of the hydrogen spectrum is .4118 x 10-4 cm. Use this value in the Rydberg equation to calculate the RH value using n1 = 2, and n2 = 4.
Answer #1

1/lambda = - R* (1/nf^2 - 1/ni^2)

Here:

lambda = 0.4118*10^-4 cm = 4.118*10^-7 m

nf = 4

ni = 2

Use:

1/(4.118*10^-7 m) = - R* (1/4^2 - 1/2^2)

1/(4.118*10^-7 m) = - R* (1/16 - 1/4)

1/(4.118*10^-7 m) = - R* ((1-4)/16)

1/(4.118*10^-7 m) = R* (3/16)

R = 16/(3*4.118*10^-7 m)

= 1.295*10^7 m-1

Answer: 1.295*10^7 m-1